IR CAVITY RINGDOWN STUDY OF HOONO FORMATION FROM OH + NO_{2}

SERGEY NIZKORODOV, BRIAN D. BEAN, ANDREW J. MOLLNER, \underline{MITCHIO OKUMURA}, Chemistry Dept. California Institute of Technology, Pasadena, CA 91125; and S. P. SANDER, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr., Pasadena, CA 91109.

The reaction of OH + NO_{2} → HNO_{3} is a critical sink of HO_{x} and NO_{x} radicals in the atmosphere. The unstable intermediate HOONO has long been proposed as a minor channel. Because HOONO is predicted to be bound by <20 kcal/mol, it would thermally decompose to reactants and would not serve as a sink for OH radicals; hence, its formation would reduce the rate of stable nitric acid formation. We report direct detection of this adduct at 3280cm^{−1}by IR cavity ringdown spectroscopy in a discharge flow cell, and present a study of its relative yield and significance in the troposphere and stratosphere.